# Consistent Story Generation: Unlocking the Potential of Zigzag Sampling

### v4

@@ 1 Introduction
In recent years, breakthroughs in visual generation techniques have fundamentally transformed the way visual content is created. State-of-the-art image [1, 2, 3, 4] and video generation models [5, 6, 7, 8, 9, 10] now enable users to produce highly diverse visual outputs with remarkable realism and flexibility. These models can be guided by various forms of control, including bounding boxes [11, 12], object motion trajectories [13, 14], image prompts [5, 4], and even brain signal inputs [15, 16, 17], significantly expanding the creative possibilities in both professional and everyday settings.
@@
Despite these impressive advances, challenges remain, particularly in visual storytelling tasks, where multiple pieces of visual content must consistently preserve the identity and key characteristics of subjects across scenes, three examples are presented in Figure 1. A dominant approach to this problem is personalization, which involves learning a subject-specific embedding  [18, 19, 20, 21, 22, 23]. While effective, this method has notable limitations: the need for per-subject fine-tuning makes it resource-intensive and difficult to scale, and it often leads to overfitting, which can degrade the model’s general understanding of text and reduce its generative diversity. To address these limitations, past research has explored tuning-free methods [24, 25, 26], which encode subject information using a unified image encoder and leverage large-scale pretraining. These methods eliminate the need for per-subject optimization and offer greater scalability. However, they rely on high-quality, large-scale personalization datasets that are costly to curate and demand substantial computational resources for effective training.
More recently, a line of training-free methods has emerged, targeting the specific challenge of subject-consistent visual story generation without fine-tuning or extensive subject-specific data. The StoryDiffusion model [27] introduces consistent attention by constructing image-level interactions through the random sharing of visual features across different frames. In contrast, the ConsistentStory model [28] aims to preserve subject identity by establishing visual interactions between images more selectively. However, while both methods show promise in maintaining subject consistency, they suffer from notable drawbacks: a reduced ability to follow textual prompts and diminished generative diversity. These issues arise from their direct interference with the diffusion model’s generative process. Departing from these strategies, the One-Prompt One-Story model [29] proposes a different approach. First, it fuses all the prompts from a story into a single extended prompt. Then, it generates each image by reweighting the prompt embeddings to reflect the desired focus for each scene. In addition, it improves text-to-image attention by reinforcing subject-relevant information within the prompt itself. Although this model achieves good prompt fidelity, it struggles to maintain subject consistency, particularly when the prompt offers limited detail about the subject.
@@
As shown in Figure 2, our method decomposes each generation step into three sub-steps—zig, zag, and generation. It proposes the combination of two key components: Zig Visual Sharing (ZVS) and Asymmetric Prompt Zigzag Inference (APZI). In the ZVS module, subject-specific visual information that is extracted using a subject-only prompt is injected into the self-attention layers during the generation of each scene. This allows the model to encode subject identity directly into the latent representations without disrupting overall scene composition. To further balance subject consistency with textual fidelity, APZI introduces an asymmetric prompt schedule: the zig and generation steps leverage the one-prompt technique proposed in [29], while the zag step employs a null prompt. This asymmetry prevents overfitting to textual descriptions and provides the model with a dedicated phase to integrate subject information. Together, ZVS and APZI enable effective subject conditioning in the latent space while preserving the model’s ability to follow textual prompts, resulting in more coherent and consistent multiscene image generation.
We evaluate our method on two representative text-to-image architectures: the U-Net-based SDXL model [30] and the transformer-based FLUX model [31]. Experimental results demonstrate that our approach significantly improves subject consistency in generated images while maintaining strong alignment with textual prompts, outperforming existing baselines. In summary, our main contributions are:
@@ 2.1 Diffusion Models
Diffusion models are a class of generative models that have demonstrated remarkable success in synthesizing high-quality images [1, 30, 4, 3], videos [9, 7, 5, 6], and 3D content [32, 33, 34]. These models consist of a forward (noising) process and a reverse (denoising) process. In the forward process, a clean image is progressively corrupted by the addition of Gaussian noise over several steps. During the reverse process, a neural network learns to gradually denoise a random Gaussian sample, ultimately reconstructing a coherent image. The multi-step nature of this inference pipeline offers significant flexibility for controllable or guided generation tasks, such as layout-constrained image synthesis [12, 35, 36, 11], motion-controlled video generation [13, 14], and subject-consistent image personalization [19, 18]. Despite their strong performance, diffusion models remain an active area of research, with ongoing efforts to expand their generative capabilities and improve efficiency. For instance, recent works [37, 38] have proposed accelerated sampling strategies that reduce the number of required inference steps to fewer than $50$. Other studies address issues such as exposure bias [39, 40], leading to more robust and higher-fidelity outputs. In general, diffusion models continue to evolve rapidly, and innovations in architecture, training strategies, and sampling techniques push the boundaries of generative modeling.
@@ 2.2 Training Based Consistent Text to Image Generation
Since the release of powerful open-source text-to-image diffusion models, generating images featuring consistent subjects — whether for image personalization or visual storytelling — has become an increasingly active area of research. Early approaches [18, 19, 41, 23, 20, 21, 22] primarily rely on test-time tuning techniques. These methods typically fine-tune either the entire diffusion network or specific components, learning a unique subject embedding or a small auxiliary network from a few reference images. Although these methods achieve high-quality results in subject-driven image generation, their reliance on test-time optimization poses significant scalability challenges. To address this limitation, a growing body of work [42, 25, 43, 44, 45, 46, 47] has explored encoder-based strategies. These methods utilize image encoders to extract subject-specific features from reference images and inject the resulting embeddings into the diffusion model to guide subject-consistent generation. By avoiding per-subject fine-tuning, such approaches improve efficiency and scalability in real-world applications.
@@ 2.3 Training-Free Consistent Text to Image Generation
In parallel with training-based approaches, training-free methods for consistent text-to-image generation have also received significant attention in the visual generation community. A popular direction involves leveraging the attention mechanism to identify subject-relevant visual features, which are then reused to guide subsequent image generation. Initially, such techniques were explored in tasks like appearance transfer [48], image editing [49, 50], and image translation [51]. More recently, ConsisStory [28] adapted this idea to visual storytelling. Specifically, it uses cross-attention scores between text and image tokens to identify subject-relevant visual features in one image, and then injects these features into the generation of subsequent images to maintain subject consistency. In contrast to attention-based guidance, the current state-of-the-art method, 1Prompt1Story [29], maintains subject consistency by composing a unified prompt that combines all scene descriptions. During generation, the model adjusts the weighting of each sub-prompt depending on the scene. Our approach differs from both strategies by introducing asymmetric guidance within the zigzag sampling framework. This asymmetric design strengthens subject consistency across generated images while preserving the model’s ability to follow textual prompts faithfully.
@@ 3.1 Preliminary
Latent Diffusion Model. The Latent Diffusion Model (LDM) comprises an autoencoder—consisting of an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$—that maps between pixel space and latent space. A diffusion model $\epsilon_{\theta}$, parameterized by $\theta$, is trained to model the noise in the latent space. For text-conditioned generation, a frozen text encoder $\tau_{\zeta}$ is used to embed the input text $\mathcal{P}$ into a dense representation [4]. The diffusion model is trained using the following loss function:
@@
During training, the diffusion process involves predicting noise at a randomly sampled time step $t$, drawn from a uniform distribution. To effectively model this denoising process—especially in the context of text-to-image generation—attention mechanisms play a central role in the architecture of diffusion models. In UNet-based models such as SDXL [30], cross-attention layers facilitate the integration of textual information by aligning latent visual features with text embeddings. This allows the model to generate images that are semantically consistent with the input text. Simultaneously, self-attention layers help capture spatial and semantic relationships within the visual latent space, enabling more coherent and detailed image synthesis. In contrast, the transformer-based FLUX model [31] adopts a different strategy. It uses a unified self-attention mechanism with modality-specific projection layers for visual and textual inputs. This design allows FLUX to fuse semantic information across modalities without relying on explicit cross-attention, leveraging the transformer’s strength in modeling complex dependencies. Our method is seamlessly integrated in the UNet-based model SDXL and the transformer-based model FLUX.
@@
To address the challenge of balancing subject consistency and text fidelity in story-based image generation, we propose a novel Asymmetric Zigzag Sampling strategy that leverages the strengths of both diffusion-based generation and semantic conditioning. Prior work has introduced zigzag sampling—a method that decomposes each diffusion denoising step into three sub-steps: zig, zag, and generation—to improve the performance of generative models. Inspired by the distinct functional roles identified in recent work [52], where the zig step facilitates exploration, the zag step enables refinement, and the generation step produces the final output, we introduce asymmetry into this process to more precisely regulate the flow of semantic information across steps. In our design, identity visual semantics are injected exclusively during the zig (exploration) step to establish strong identity grounding early in the denoising trajectory. The zag step then adjusts the latent space without further visual input, helping to avoid overfitting to the identity. Finally, the generation step focuses solely on prompt alignment, free from additional visual conditioning. This asymmetric configuration allows our method to achieve a superior balance between subject consistency and prompt fidelity, particularly in complex, multi-image story generation tasks. An overview of the proposed method is illustrated in Figure 3, with detailed step-by-step operations described below. Formally, let $x_{t}$ denote the noisy latent at timestep $t$, and $x_{t-1}$ the denoised latent. While the zig and generation steps follow the standard forward denoising trajectory, the zag step performs an inverse denoising operation, mapping $x_{t}$ to a prior latent $x_{t+1}$. The computation for each sub-step is governed by the noise schedule $\alpha_{t}$ predefined by the underlying diffusion model.
@@
Asymmetry Visual Guidance. As described above, we inject strong subject-specific cues into the latent representation during the zig step. To extract these cues, we first run the diffusion denoising process using an identity-focused prompt that describes only the visual characteristics of the subject. Following prior work [28], we compute text-image attention scores to identify subject-relevant visual tokens at each layer and timestep. These tokens serve as the basis for semantic injection, helping the model preserve subject identity across generated images. We denote the subject-related attention scores in the text-image attention layer as:
@@
Asymmetry Prompt Guidance. For text-based guidance, we have first adapted an approach proposed in [29], which concatenates all scene-level prompts into a single sentence and dynamically reweights the contribution of each sub-sentence during the generation of the corresponding image. This allows the model to maintain narrative coherence across scenes while emphasizing the relevant textual content at each generation step. In addition, we have integrated the Identity-Preserving Contrastive Alignment (IPCA) technique [29] to further strengthen subject representation through the text prompt, reinforcing the model’s ability to retain subject identity throughout the story. However, directly applying these text-guidance strategies within the standard zigzag sampling framework has led to suboptimal results. We hypothesize that the strong text-driven supervision during the zag step conflicts with the subject-specific information introduced through visual token injection, potentially overriding or distorting it. This interference degrades subject consistency and overall image quality. To address this issue, we introduce a novel asymmetric zigzag prompt guidance strategy. Specifically, we apply the enhanced text-guidance mechanisms only during the zig and generation steps, where they align well with semantic injection and image refinement. During the zag step, we instead use a null prompt to prevent conflicting and allow the subject-specific visual information to propagate unimpeded. This asymmetry helps preserve both textual relevance and subject consistency across the generated sequence.
@@ 4 Experiments
Comparison Methods. We integrate our method in two backbone architectures: the UNet-based SDXL [30] and the transformer-based FLUX [31] models. For the SDXL backbone, we compare our approach against several state-of-the-art baselines, including both training-based methods—The Chosen One [53], PhotoMaker [43], and IP-Adapter [24]—and training-free methods—ConsiStory [28], StoryDiffusion [27], and 1Prompt1Story [29]. For the FLUX model, as no prior methods have been demonstrated to support this architecture, we re-implement the most recent and competitive training-free method, 1Prompt1Story, on top of FLUX. We then use this as a baseline for comparison against our method. This dual-platform evaluation demonstrates the adaptability and effectiveness of our approach across diverse diffusion architectures.
Evaluation Metrics. Following prior work [29, 28, 27], we evaluate all models along two key dimensions: prompt alignment and subject consistency. To assess prompt alignment, we use the CLIP image and text encoders to compute the average CLIP-Score [54] between each generated image and its corresponding prompt. This metric reflects how well the visual content of the generated image aligns with the intended textual description. For subject consistency, we adopt two complementary metrics from previous studies [29]: DreamSim [55] and CLIP-I [54]. DreamSim measures perceptual similarity and has shown strong correlation with human judgment in evaluating visual coherence. CLIP-I computes the average cosine similarity between CLIP image embeddings, capturing the consistency of subject identity across different images. To ensure that subject consistency is not influenced by variations in background content, we follow prior evaluation protocols and apply CarveKit [56] to remove the background from each generated image. The removed regions are then filled with random noise, isolating the subject and allowing for a more focused and accurate evaluation.

### v5

@@ 1 Introduction
In recent years, breakthroughs in visual generation techniques have fundamentally transformed the way visual content is created. State-of-the-art image [sd3, dalle-e, pixart, stable-diffusion] and video generation models [hunyuanvideo, wanxiang, step-video, magi1, vidu, moviegen] now enable users to produce highly diverse visual outputs with remarkable realism and flexibility. These models can be guided by various forms of control, including bounding boxes [gligen, layoutdiffusion], object motion trajectories [animate-motion, motion-prompt], image prompts [hunyuanvideo, stable-diffusion], and even brain signal inputs [sun2023contrast, sun2023decoding, sun2025neuralflix], significantly expanding the creative possibilities in both professional and everyday settings.
@@
Despite these impressive advances, challenges remain, particularly in visual storytelling tasks, where multiple pieces of visual content must consistently preserve the identity and key characteristics of subjects across scenes, three examples are presented in Figure 1. A dominant approach to this problem is personalization, which involves learning a subject-specific embedding  [dreambooth, text_inversion, customize, key-lock, neural-space, towards]. While effective, this method has notable limitations: the need for per-subject fine-tuning makes it resource-intensive and difficult to scale, and it often leads to overfitting, which can degrade the model’s general understanding of text and reduce its generative diversity. To address these limitations, past research has explored tuning-free methods [ip-adapter, id-adapter, blip-dif], which encode subject information using a unified image encoder and leverage large-scale pretraining. These methods eliminate the need for per-subject optimization and offer greater scalability. However, they rely on high-quality, large-scale personalization datasets that are costly to curate and demand substantial computational resources for effective training.
More recently, a line of training-free methods has emerged, targeting the specific challenge of subject-consistent visual story generation without fine-tuning or extensive subject-specific data. The StoryDiffusion model [storydiffusion] introduces consistent attention by constructing image-level interactions through the random sharing of visual features across different frames. In contrast, the ConsistentStory model [consistent-story] aims to preserve subject identity by establishing visual interactions between images more selectively. However, while both methods show promise in maintaining subject consistency, they suffer from notable drawbacks: a reduced ability to follow textual prompts and diminished generative diversity. These issues arise from their direct interference with the diffusion model’s generative process. Departing from these strategies, the One-Prompt One-Story model [one-prompt] proposes a different approach. First, it fuses all the prompts from a story into a single extended prompt. Then, it generates each image by reweighting the prompt embeddings to reflect the desired focus for each scene. In addition, it improves text-to-image attention by reinforcing subject-relevant information within the prompt itself. Although this model achieves good prompt fidelity, it struggles to maintain subject consistency, particularly when the prompt offers limited detail about the subject.
@@
As shown in Figure 2, our method decomposes each generation step into three sub-steps—zig, zag, and generation. It proposes the combination of two key components: Zig Visual Sharing (ZVS) and Asymmetric Prompt Zigzag Inference (APZI). In the ZVS module, subject-specific visual information that is extracted using a subject-only prompt is injected into the self-attention layers during the generation of each scene. This allows the model to encode subject identity directly into the latent representations without disrupting overall scene composition. To further balance subject consistency with textual fidelity, APZI introduces an asymmetric prompt schedule: the zig and generation steps leverage the one-prompt technique proposed in [one-prompt], while the zag step employs a null prompt. This asymmetry prevents overfitting to textual descriptions and provides the model with a dedicated phase to integrate subject information. Together, ZVS and APZI enable effective subject conditioning in the latent space while preserving the model’s ability to follow textual prompts, resulting in more coherent and consistent multiscene image generation.
We evaluate our method on two representative text-to-image architectures: the U-Net-based SDXL model [sdxl] and the transformer-based FLUX model [flux]. Experimental results demonstrate that our approach significantly improves subject consistency in generated images while maintaining strong alignment with textual prompts, outperforming existing baselines. In summary, our main contributions are:
@@ 2.1 Diffusion Models
Diffusion models are a class of generative models that have demonstrated remarkable success in synthesizing high-quality images [sd3, sdxl, stable-diffusion, pixart], videos [vidu, step-video, hunyuanvideo, wanxiang], and 3D content [magic3d, dreamfusion, stable3d]. These models consist of a forward (noising) process and a reverse (denoising) process. In the forward process, a clean image is progressively corrupted by the addition of Gaussian noise over several steps. During the reverse process, a neural network learns to gradually denoise a random Gaussian sample, ultimately reconstructing a coherent image. The multi-step nature of this inference pipeline offers significant flexibility for controllable or guided generation tasks, such as layout-constrained image synthesis [layoutdiffusion, t2i, adding, gligen], motion-controlled video generation [animate-motion, motion-prompt], and subject-consistent image personalization [text_inversion, dreambooth]. Despite their strong performance, diffusion models remain an active area of research, with ongoing efforts to expand their generative capabilities and improve efficiency. For instance, recent works [denoising, dpm] have proposed accelerated sampling strategies that reduce the number of required inference steps to fewer than $50$. Other studies address issues such as exposure bias [alleviating, elucidating], leading to more robust and higher-fidelity outputs. In general, diffusion models continue to evolve rapidly, and innovations in architecture, training strategies, and sampling techniques push the boundaries of generative modeling.
@@ 2.2 Training Based Consistent Text to Image Generation
Since the release of powerful open-source text-to-image diffusion models, generating images featuring consistent subjects — whether for image personalization or visual storytelling — has become an increasingly active area of research. Early approaches [dreambooth, text_inversion, svdiff, towards, customize, key-lock, neural-space] primarily rely on test-time tuning techniques. These methods typically fine-tune either the entire diffusion network or specific components, learning a unique subject embedding or a small auxiliary network from a few reference images. Although these methods achieve high-quality results in subject-driven image generation, their reliance on test-time optimization poses significant scalability challenges. To address this limitation, a growing body of work [blip-diffusion, id-adapter, photomaker, storymaker, visualpersona, elite, subject] has explored encoder-based strategies. These methods utilize image encoders to extract subject-specific features from reference images and inject the resulting embeddings into the diffusion model to guide subject-consistent generation. By avoiding per-subject fine-tuning, such approaches improve efficiency and scalability in real-world applications.
@@ 2.3 Training-Free Consistent Text to Image Generation
In parallel with training-based approaches, training-free methods for consistent text-to-image generation have also received significant attention in the visual generation community. A popular direction involves leveraging the attention mechanism to identify subject-relevant visual features, which are then reused to guide subsequent image generation. Initially, such techniques were explored in tasks like appearance transfer [cross-transfer], image editing [hertz2022prompt, plug], and image translation [image_translate]. More recently, ConsisStory [consistent-story] adapted this idea to visual storytelling. Specifically, it uses cross-attention scores between text and image tokens to identify subject-relevant visual features in one image, and then injects these features into the generation of subsequent images to maintain subject consistency. In contrast to attention-based guidance, the current state-of-the-art method, 1Prompt1Story [one-prompt], maintains subject consistency by composing a unified prompt that combines all scene descriptions. During generation, the model adjusts the weighting of each sub-prompt depending on the scene. Our approach differs from both strategies by introducing asymmetric guidance within the zigzag sampling framework. This asymmetric design strengthens subject consistency across generated images while preserving the model’s ability to follow textual prompts faithfully.
@@ 3.1 Preliminary
Latent Diffusion Model. The Latent Diffusion Model (LDM) comprises an autoencoder—consisting of an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$—that maps between pixel space and latent space. A diffusion model $\epsilon_{\theta}$, parameterized by $\theta$, is trained to model the noise in the latent space. For text-conditioned generation, a frozen text encoder $\tau_{\zeta}$ is used to embed the input text $\mathcal{P}$ into a dense representation [stable-diffusion]. The diffusion model is trained using the following loss function:
@@
During training, the diffusion process involves predicting noise at a randomly sampled time step $t$, drawn from a uniform distribution. To effectively model this denoising process—especially in the context of text-to-image generation—attention mechanisms play a central role in the architecture of diffusion models. In UNet-based models such as SDXL [sdxl], cross-attention layers facilitate the integration of textual information by aligning latent visual features with text embeddings. This allows the model to generate images that are semantically consistent with the input text. Simultaneously, self-attention layers help capture spatial and semantic relationships within the visual latent space, enabling more coherent and detailed image synthesis. In contrast, the transformer-based FLUX model [flux] adopts a different strategy. It uses a unified self-attention mechanism with modality-specific projection layers for visual and textual inputs. This design allows FLUX to fuse semantic information across modalities without relying on explicit cross-attention, leveraging the transformer’s strength in modeling complex dependencies. Our method is seamlessly integrated in the UNet-based model SDXL and the transformer-based model FLUX.
@@
To address the challenge of balancing subject consistency and text fidelity in story-based image generation, we propose a novel Asymmetric Zigzag Sampling strategy that leverages the strengths of both diffusion-based generation and semantic conditioning. Prior work has introduced zigzag sampling [bai2024zigzag, bai2025weak]—a method that decomposes each diffusion denoising step into three sub-steps: zig, zag, and generation—to improve the performance of generative models. Inspired by the distinct functional roles identified in recent work [bai2024zigzag, bai2025weak], where the zig step facilitates exploration, the zag step enables refinement, and the generation step produces the final output, we introduce asymmetry into this process to more precisely regulate the flow of semantic information across steps. In our design, identity visual semantics are injected exclusively during the zig (exploration) step to establish strong identity grounding early in the denoising trajectory. The zag step then adjusts the latent space without further visual input, helping to avoid overfitting to the identity. Finally, the generation step focuses solely on prompt alignment, free from additional visual conditioning. This asymmetric configuration allows our method to achieve a superior balance between subject consistency and prompt fidelity, particularly in complex, multi-image story generation tasks. An overview of the proposed method is illustrated in Figure 3, with detailed step-by-step operations described below. Formally, let $x_{t}$ denote the noisy latent at timestep $t$, and $x_{t-1}$ the denoised latent. While the zig and generation steps follow the standard forward denoising trajectory, the zag step performs an inverse denoising operation, mapping $x_{t}$ to a prior latent $x_{t+1}$. The computation for each sub-step is governed by the noise schedule $\alpha_{t}$ predefined by the underlying diffusion model.
@@
Asymmetry Visual Guidance. As described above, we inject strong subject-specific cues into the latent representation during the zig step. To extract these cues, we first run the diffusion denoising process using an identity-focused prompt that describes only the visual characteristics of the subject. Following prior work [consistent-story], we compute text-image attention scores to identify subject-relevant visual tokens at each layer and timestep. These tokens serve as the basis for semantic injection, helping the model preserve subject identity across generated images. We denote the subject-related attention scores in the text-image attention layer as:
@@
Asymmetry Prompt Guidance. For text-based guidance, we have first adapted an approach proposed in [one-prompt], which concatenates all scene-level prompts into a single sentence and dynamically reweights the contribution of each sub-sentence during the generation of the corresponding image. This allows the model to maintain narrative coherence across scenes while emphasizing the relevant textual content at each generation step. In addition, we have integrated the Identity-Preserving Contrastive Alignment (IPCA) technique [one-prompt] to further strengthen subject representation through the text prompt, reinforcing the model’s ability to retain subject identity throughout the story. However, directly applying these text-guidance strategies within the standard zigzag sampling framework has led to suboptimal results. We hypothesize that the strong text-driven supervision during the zag step conflicts with the subject-specific information introduced through visual token injection, potentially overriding or distorting it. This interference degrades subject consistency and overall image quality. To address this issue, we introduce a novel asymmetric zigzag prompt guidance strategy. Specifically, we apply the enhanced text-guidance mechanisms only during the zig and generation steps, where they align well with semantic injection and image refinement. During the zag step, we instead use a null prompt to prevent conflicting and allow the subject-specific visual information to propagate unimpeded. This asymmetry helps preserve both textual relevance and subject consistency across the generated sequence.
@@ 4 Experiments
Comparison Methods. We integrate our method in two backbone architectures: the UNet-based SDXL [sdxl] and the transformer-based FLUX [flux] models. For the SDXL backbone, we compare our approach against several state-of-the-art baselines, including both training-based methods—The Chosen One [chosen-one], PhotoMaker [photomaker], and IP-Adapter [ip-adapter]—and training-free methods—ConsiStory [consistent-story], StoryDiffusion [storydiffusion], and 1Prompt1Story [one-prompt]. For the FLUX model, as no prior methods have been demonstrated to support this architecture, we re-implement the most recent and competitive training-free method, 1Prompt1Story, on top of FLUX. We then use this as a baseline for comparison against our method. This dual-platform evaluation demonstrates the adaptability and effectiveness of our approach across diverse diffusion architectures.
Evaluation Metrics. Following prior work [one-prompt, consistent-story, storydiffusion], we evaluate all models along two key dimensions: prompt alignment and subject consistency. To assess prompt alignment, we use the CLIP image and text encoders to compute the average CLIP-Score [clipscore] between each generated image and its corresponding prompt. This metric reflects how well the visual content of the generated image aligns with the intended textual description. For subject consistency, we adopt two complementary metrics from previous studies [one-prompt]: DreamSim [dreamsim] and CLIP-I [clipscore]. DreamSim measures perceptual similarity and has shown strong correlation with human judgment in evaluating visual coherence. CLIP-I computes the average cosine similarity between CLIP image embeddings, capturing the consistency of subject identity across different images. To ensure that subject consistency is not influenced by variations in background content, we follow prior evaluation protocols and apply CarveKit [carvekit] to remove the background from each generated image. The removed regions are then filled with random noise, isolating the subject and allowing for a more focused and accurate evaluation.